# High Performance Space Debris Tracking in Complex Skylight Backgrounds with a Large-Scale Dataset

### v2

@@ 5.4 Ablation Study
To analyze the contributions of different components in SDT-Net, we conduct an ablation study on the Debris Test and Dense Debris Test sets, as shown in Tab 4. We evaluate three configurations: (a) Centertrack, which predicts line source debris using center points and box size parameters; (b) LSE (Line Source Embedding), which models debris as line sources and directly locates their endpoint pairs; and (c) RoI-FE (Region-of-Interest Feature Enhancement), which further enhances debris feature representation.
@@
Region-of-Interest Feature Enhancement (RoI-FE): Further incorporating segmentation masks (row c) enhances the representation of debris features. This results in additional performance gains, with MOTA improving from 85.5% to 87.7% on the Debris Test set and from 64.9% to 70.3% on the Dense Debris Test set. Similarly, HOTA sees a significant increase, confirming that segmentation helps improve tracking robustness.
@@
LSE RoI-FE Debris Test Dense Debris Test
IDF1↑ MOTA↑ HOTA↑ IDF1↑ MOTA↑ HOTA↑
@@
Table 4: Evaluation the Antarctic station data on SDT-Net.
Method MOTA HOTA Video Frame

### v3

@@ 5.4 Ablation Study
To analyze the contributions of different components in SDT-Net, we conduct an ablation study on the Debris Test and Dense Debris Test sets, as shown in Tab 3. We evaluate three configurations: (a) Based on Centertrack, the center point and box size parameters are used to predict line source debris, and the debris offset module is also included; (b) LSE (Line Source Embedding), which models debris as line sources and directly locates their endpoint pairs; (c) RoI-FE (Region of Interest Feature Enhancement) further enhances the fragment feature representation and removes the debris offset module; and (d) Complete SDT-Net.
@@
Region-of-Interest Feature Enhancement (RoI-FE): Further incorporating segmentation masks (compare rows b and d) enhances the representation of debris features. This results in additional performance gains, with MOTA improving from 85.5% to 87.7% on the Debris Test set and from 64.9% to 70.3% on the Dense Debris Test set. Similarly, HOTA sees a significant increase, confirming that segmentation helps improve tracking robustness.
Offset Module: As shown in rows (c) and (d), introducing the debris offset module alone can significantly improve all indicators. Specifically, MOTA is improved from 86.0% (row c) to 87.7% on the fragmented test set, and from 62.9% to 70.3% on the dense fragmented test set. These improvements show that the offset module plays a key role in improving endpoint accuracy and overall tracking robustness. Finally, after integrating all modules, IDF1 is 80.6%, MOTA is 70.3%, and HOTA is 73.6% on the dense fragmented test set.
@@
LSE RoI-FE Offset Module Debris Test Dense Debris Test
[…]
@@
Table 4: Evaluation of Antarctic station data on SDT-Net.
[…]